OpenStudy (anonymous):

One number is four more than a second number. Two times the first number is 12 more than four times the second number. a. - 10 and - 14 b. 1 and - 3 c. 2 and - 2 d. 3 and - 1 One number is four more than a second number. Two times the first number is 12 more than four times the second number. a. - 10 and - 14 b. 1 and - 3 c. 2 and - 2 d. 3 and - 1 @Mathematics

OpenStudy (anonymous):

Let us start peeling the problem step at a time. Anytime you have a problem talking about some unknown numbers, assume they are x and y. In this case, let us say the first number is x and second number is y. We don't know what they are yet. Then, transform words into equations. One number is four more than the second number => x = y + 4 (four more) Two times the first number is 12 more than four times the second number => 2x = 12 + 4y. You need now solve these two equations: x = y + 4 2x = 4y + 12

OpenStudy (anonymous):

i got D?

OpenStudy (anonymous):

2(y+4) = 4y + 12 => 2y + 8 = 4y + 12 => 2y -4y = 12 -8 => -2y = 4 => y = -2 => x = -2 + 4 => x = 2

OpenStudy (anonymous):

You can verify if your answer is right or wrong by plugging in the values of x and y you got into the original two equations and see if they compute.

OpenStudy (anonymous):

Thank you so much!